OpenStudy (anonymous):

y varies jointly with p and g. when p=3 and q=5, y=60. find y when p=2 and q=6

OpenStudy (kropot72):

y=kpq ..........(1) where k is a constant of proportionality. We need to find the value of k by plugging the given values of p, q and y into equation (1) and solving for k. 60 = k * 3 * 5 60 = 15 * k ...........(2) Can you solve equation (2) to find the value of k? When you have the value of k, values can be plugged into equation (1) to find the solution to the question.